### v1

# ModelNet40-E: An Uncertainty-Aware Benchmark for Point Cloud Classification

Abstract
We introduce ModelNet40-E, a new benchmark designed to assess the robustness and calibration of point cloud classification models under synthetic LiDAR-like noise. Unlike existing benchmarks, ModelNet40-E provides both noise-corrupted point clouds and point-wise uncertainty annotations via Gaussian noise parameters ($\sigma$, $\mu$), enabling fine-grained evaluation of uncertainty modeling. We evaluate three popular models—PointNet, DGCNN, and Point Transformer v3—across multiple noise levels using classification accuracy, calibration metrics, and uncertainty-awareness. While all models degrade under increasing noise, Point Transformer v3 demonstrates superior calibration, with predicted uncertainties more closely aligned with the underlying measurement uncertainty.
Introduction
In recent years, 3D point cloud classification has gained increasing importance across a wide range of applications, including autonomous driving (Lang et al. 2018; Yan, Mao, and Li 2018; Sun et al. 2019), robotics (Mousavian, Eppner, and Fox 2019), augmented reality (Izadi et al. 2011; Newcombe, Fox, and Seitz 2015; Jiang et al. 2025; Newcombe et al. 2011; Yu et al. 2020), and remote sensing (Wang et al. 2025, 2024; Pan et al. 2025). In response, numerous models have been proposed, ranging from early architectures like PointNet (Qi et al. 2017a) and PointNet++ (Qi et al. 2017b) to more advanced designs such as DGCNN (Wang et al. 2019), Point2Vec (Abou Zeid et al. 2023), and Point Transformer (Zhao et al. 2021; Wu et al. 2022, 2024).
Constrained by the limited availability of real-world annotated point cloud data, most of these models are trained and evaluated on synthetic datasets, such as ModelNet40 (Wu et al. 2015) and Shapenet (Chang et al. 2015). These datasets are composed of clean, idealized point clouds that lack the noise, occlusions, and distortions typically present in real sensor data. As a result, models that perform well in these benchmarks often fail to generalize to real-world environments (Wu et al. 2018; Zhao, Zhao, and Lee 2022; Ren, Yang, and Velipasalar 2025).
To ensure fair comparisons and improve the applicability of the models in real-world scenarios, robustness benchmarks like ModelNet-C (Sun et al. 2022) have been proposed. These benchmarks introduce synthetic corruptions such as rotations, jittering, scaling, and point dropout to test the resilience of models to input perturbations. Despite offering valuable insights, they face key limitations: (1) they apply generic perturbations that do not reflect realistic LiDAR noise, and (2) they lack ground-truth uncertainty annotations. In safety-critical applications, such as autonomous driving and robotic manipulation, it is not enough for a model to simply output a class label. Equally important is assessing how confident the model is in its predictions. Overconfident yet incorrect predictions can lead to catastrophic failures, making calibration and uncertainty awareness essential attributes for building reliable models.
We introduce ModelNet40-E, a novel benchmark that builds upon the original ModelNet40 dataset by corrupting it with LiDAR-like noise at three severity levels: Light, Moderate, and Heavy. Unlike random jitter or dropout, LiDAR noise is range-dependent, anisotropic, and often structured, influenced by surface reflectivity, incidence angles, and environmental conditions. Capturing these complexities is critical for creating realistic benchmarks that truly test model robustness and uncertainty estimation. Our noise simulation is inspired by empirical studies of LiDAR measurement errors (Gschwandtner et al. 2011), incorporating range-dependent Gaussian noise, incidence angle effects, systematic bias, and random outliers. Crucially, for every point in the dataset, we also provide the corresponding per-point noise statistics: standard deviation $\sigma$ and expected bias $\mu$. This allows for rigorous evaluation of both predictive performance and uncertainty estimation.
The main contributions of this work are as follows:
  • We introduce ModelNet40-E, a realistic, uncertainty-aware benchmark for point cloud classification.
  • We evaluate three popular point cloud models—PointNet, DGCNN, and PTv3—across several metrics for different noise levels.
  • We show that while all models degrade under noise, PTv3 demonstrates superior robustness and calibration, particularly in aligning predicted confidence with ground-truth uncertainty.
Related Work
Point Cloud Classification Benchmarks
[…]
However, such synthetic datasets do not accurately reflect real-world sensor data, which is often affected by noise, occlusions, and incomplete scans. Consequently, models trained solely on synthetic data often struggle to generalize to real-world scenarios.
To address this limitation, Sun et al. (2022) proposed ModelNet40-C, a corrupted version of ModelNet40 that introduces a wide range of synthetic perturbations, including Gaussian noise, scaling, rotation, translation, point dropout, and jitter. These corruptions are designed to simulate various distortions that models might encounter in real-world data. Sun et al. (2022) evaluate robustness by measuring classification accuracy under each corruption type and severity level.
Robustness and Adversarial Attacks on Point Clouds
Robustness in point cloud classification refers to a model’s ability to maintain performance under various perturbations. These include sensor noise, occlusions, point sparsity, and misalignment, which are common challenges present in real-world data. To evaluate resilience under such conditions, recent benchmarks like ModelNet40-C (Sun et al. 2022) apply a wide range of synthetic corruptions (e.g., rotation, scaling, dropout, jitter) and measure performance across different perturbation levels.
Beyond natural corruptions, there has been growing interest in studying adversarial robustness, where small, carefully crafted perturbations are designed to fool classifiers. Several works have extended adversarial attack strategies from the 2D image domain to 3D point clouds. For instance, Xiang, Qi, and Li (2018) proposes adversarial perturbation and point generation strategies to mislead PointNet into high-confidence misclassifications. In addition, Zheng et al. (2018) introduces a point perturbation attack that uses gradient information to identify and remove critical points, degrading classification accuracy. Other approaches, like Liu, Yu, and Su (2019) and Tsai et al. (2020), have explored transformations and translations to mislead classifiers while preserving overall object shape.
To counter these attacks, a variety of defense strategies have been developed, including adversarial training, which improves model robustness by exposing it to adversarial examples during training (Sun et al. 2020), as well as input reconstruction methods that denoise or repair corrupted point clouds before classification (Zhou et al. 2018). However, many of these defenses come at the cost of reduced clean accuracy, and few have been evaluated under realistic sensor noise conditions, limiting their practical effectiveness.
While adversarial attacks expose important model vulnerabilities, robustness to natural corruptions is arguably more critical for real-world deployment. In practical applications such as autonomous driving, models must remain reliable in the presence of LiDAR noise, occlusions, and imperfect sensor inputs. However, most existing benchmarks focus on synthetic noise, often failing to capture the complexity of real-world 3D sensing environments.
Uncertainty Estimation in Deep Learning
Uncertainty estimation in deep learning refers to a model’s ability to not only make predictions but also to quantify how confident it is in those predictions. In a well-calibrated model, confidence scores should decrease appropriately as input perturbations are introduced. If confidence scores remain high even in the presence of significant corruptions, this may lead to overconfident misclassifications. It is well established that neural networks are often miscalibrated and tend to be overconfident, particularly as model capacity or training time increases (Guo et al. 2017).
Uncertainty is often categorized into two types: aleatoric and epistemic. Aleatoric uncertainty refers to the inherent noise in the data, such as sensor measurement errors or occlusions in point clouds. This type of uncertainty cannot be reduced by introducing more data. In contrast, epistemic uncertainty arises from the model’s ignorance due to limited or imbalanced data and can be reduced by improving data diversity or quantity.
Several techniques exist to estimate these uncertainties. For example, Gal and Ghahramani (2016) proposes the Monte Carlo (MC) dropout method, which performs multiple stochastic forward passes with dropout enabled at test time to approximate predictive uncertainty. In addition, deep ensembles (Lakshminarayanan, Pritzel, and Blundell 2017) train multiple models with different initial conditions and estimate uncertainty by measuring the variance between their outputs.
ModelNet40-E
Benchmark Overview
ModelNet40-E is an uncertainty-aware benchmark for point cloud classification. It is built upon the original ModelNet40 dataset (Wu et al. 2015), introducing several levels of LiDAR-like noise based on realistic sensor noise. By providing per-point noise statistics ($\sigma$ and $\mu$), the main goal of ModelNet40-E is to serve as a common benchmark for evaluating the uncertainty estimation capabilities of point cloud classification models.
We introduce three different levels of noise to the original ModelNet40 dataset: Light, Moderate, and Heavy. The number of samples and categories are identical to the original dataset, as well as the training and test splits.
LiDAR Noise Simulation
To approximate realistic sensor noise, we simulate a LiDAR acquisition process inspired by empirical studies of LiDAR measurement errors (Gschwandtner et al. 2011). For each clean point cloud, we compute per-point range and incidence angle relative to the sensor. Measurement noise is sampled from a Gaussian distribution whose standard deviation increases with distance and varies systematically with surface orientation. Specifically, the per-point noise $\epsilon$ is sampled as
$\epsilon\sim\mathcal{N}\bigl{(}\mu(\theta),\sigma(r,\theta)\bigr{)},$
where $r$ denotes the distance to the sensor, and $\theta$ is the incidence angle between the incoming ray and the estimated local surface normal. In particular, the noise model includes the following parameters:
Range-dependent noise. LiDAR noise increases with range. To model this, we define the range-dependent noise as
$\sigma_{range}=a+b\cdot r,$
where $r$ is the distance from the sensor to the point, $a$ is a constant base noise term (in meters), and $b$ scales noise linearly with distance.
Angle-dependent noise. Noise in LiDAR data also depends on the angle at which the laser hits the surface. The more obliquely the angle, the greater the noise. We model this effect as
$\sigma_{angle}=1+c\cdot(1-\cos{\theta}),$
where $\theta$ is the incidence angle between the ray and the surface normal, and $c$ is the scaling factor. Higher $c$ means greater sensitivity to the incidence angle.
Bias ($k$). In addition to the range-dependent noise arising from increasing measurement uncertainty with distance, there is another contribution from systematic sensor calibration offset. This effect also depends on range and is defined as
$\mu=k\cdot(1-\cos{\theta}),$
where $k$ is a deterministic bias term added to the measured range.
Outlier probability. Outliers are common in LiDAR data, often caused by spurious reflections, multipath returns, or transient measurement artifacts. We define an outlier probability parameter specifying the proportion of points that are replaced with random points—sampled uniformly within a bounding box around the object—during ModelNet40-E generation.
This procedure yields perturbed point clouds along with the ground-truth noise statistics (mean bias $\mu$ and standard deviation $\sigma$) per point.
Table 1: Configuration parameters for the light, moderate, and heavy noise levels in ModelNet40-E.
Tier Range Noise $(a,b)$ Angle Factor $c$ Bias $k$ Outlier probability
Light (0.003, 0.001) 1.5 0.005 0.01
Moderate (0.005, 0.002) 2.0 0.010 0.02
Heavy (0.010, 0.003) 3.0 0.015 0.05
Figure 1: Examples of ModelNet40 point clouds under different noise levels: none, light, moderate, and heavy. ”none” refers to the original, unperturbed point cloud.
Experiments
Experimental Setup
We evaluate three representative architectures—PointNet, DGCNN, and PTv3—on ModelNet40-E. Each model is trained independently on each noise level using the training set and evaluated on the corresponding test set. We report classification accuracy, calibration error, and uncertainty correlation.
Models
(1) PointNet, introduced by Qi et al. (2017a), was the first end-to-end architecture to directly process raw point clouds without requiring meshing or voxelization. It follows the next steps: (1) per-point feature extraction via shared multilayer perceptrons (MLPs), and (2) a max-pooling operation that aggregates the per-point features into a single global feature vector. This design allows the model to process unordered point sets and capture global shape information.
For classification tasks, the global feature vector is passed through a series of three fully connected (FC) layers that form the classification head. Each FC layer is followed by a ReLU activation and dropout, except for the last layer.
(2) Dynamic Graph CNN (DGCNN; Wang et al. 2019) is a point cloud classification and segmentation model that captures local geometric features through a key operation called EdgeConv. It starts by constructing a graph for each point based on its k-nearest neighbors (kNN) in feature space. For each edge, a local feature is computed by applying a multilayer perceptron (MLP) to the concatenation of the point’s feature and the relative offset between the point and its neighbors.
A distinctive feature of DGCNN is its dynamic graph construction: after each layer, the neighborhood graph is updated according to the new point features rather than their coordinates.
Finally, the model aggregates features across layers and uses a global pooling followed by fully connected (FC) layers to perform classification or segmentation.
(3) Point Transformer v3 (PTv3; Wu et al. 2024) is a transformer-based architecture designed for point cloud understanding. The original Point Transformer (PT; Zhao et al. 2021) first introduced attention mechanisms to capture local geometric structures in point clouds, and Point Transformer v2 (PTv2; Wu et al. 2022) further improved scalability and efficiency. PTv3 builds on these ideas, operating directly on unordered point clouds with multiple stages of attention blocks that capture both local and global dependencies. Using learned attention weights, each block aggregates contextual information by attending to nearby points.
In our implementation, we use PTv3 with only the $(x,y,z)$ coordinates as input and FlashAttention disabled. The network consists of five stages of transformer blocks followed by a max-pooling operation. A simple classification head, composed of two fully connected layers, outputs the predicted class probabilities for each sample.
Noise Configurations
We employ four different noise configurations: None, Light, Moderate, and Heavy. ”None” refers to the original, unperturbed ModelNet40 dataset. The parameters used to generate the noise are summarized in Table 1. For this benchmark, the virtual sensor position is fixed at (0,2,0)(0,-2,0)( 0 , - 2 , 0 ), which defines the range and incidence angles used in the noise simulation.
Figure 1 illustrates examples of point clouds for each noise level.
Training Details
Models are trained independently on each noise configuration (None, Light, Moderate, Heavy) for 100 epochs using cross-entropy loss as the objective. Training is performed with the Adam optimizer (Kingma 2014), using an initial learning rate of $1\times 10^{-3}$, which is decayed by a factor of 0.7 every 20 epochs. PointNet and DGCNN models are trained with a batch size of 32, while PTv3 uses a reduced batch size of 8 due to its higher memory requirement. The best models are selected based on the highest classification accuracy achieved on the test set. All experiments were conducted on a single NVIDIA RTX 4090 GPU.
Evaluation Metrics
[…]
(1) Classification Accuracy. This metric evaluates the predictive performance of classification models. It is defined as the proportion of correctly predicted test samples. In our experiments, we report accuracy in the test set for each noise level independently. It is expected that accuracy decreases as the noise level increases, and a lower decrement signals better robustness to noise.
[…]
In practice, we compute ECE by dividing the predictions into $M$ bins and comparing the average confidence and accuracy within each bin:
$\text{ECE}=\sum_{i=1}^{M}\frac{n_{i}}{N}\left|\text{acc}_{i}-\text{conf}_{i}\right|,$
where $n_{i}$ is the number of predictions in the $i^{\text{th}}$ bin, and $N$ is the total number of predictions. $\text{acc}_{i}$ and $\text{conf}_{i}$ represent the average accuracy and confidence in the $i^{\text{th}}$ bin, respectively.
(3) Correlation Between Measurement Uncertainty and Predicted Uncertainty. We evaluate whether models can accurately recognize the reliability of their predictions. This is an essential property for avoiding overconfident or underconfident outputs. To assess this, we compute the Pearson correlation between the ground truth measurement uncertainty of each sample and the model’s predicted uncertainty.
For each sample, we define the predicted uncertainty as one minus the predicted confidence, $1-p_{\text{max}}$, where $p_{\text{max}}$ is the maximum softmax probability assigned to any class. Ground truth measurement uncertainty is quantified by the mean standard deviation $\sigma$ of the simulated noise applied to that sample.
A higher Pearson correlation indicates that the model’s confidence decreases as the input noise increases, demonstrating better awareness of input uncertainty.
Main Results
[…]
Classification Accuracy
Table 2 presents the classification accuracy on the test set for all models across different noise levels. As expected, performance degrades as the noise level increases. However, this decline is notably more pronounced in PointNet compared to DGCNN and PTv3, indicating that PointNet is less robust to corruption.
Table 2: Classification Accuracy (%) for all models across different noise levels.
Model None Light Moderate Heavy
PointNet 89.30 86.30 86.18 84.64
DGCNN 91.00 89.67 89.34 89.02
PTv3 88.82 88.74 88.90 88.05
Expected Calibration Error (ECE)
Figure 4: Accuracy (left), expected calibration error (center), and Pearson correlation between predicted and measurement uncertainty (right) for PTv3 across training epochs under varying noise levels.
We evaluate the Expected Calibration Error (ECE) of all models under varying noise levels, following the procedure described in Section 2.
The results, summarized in Table 3, show that calibration error increases consistently with noise severity. This indicates that higher measurement uncertainty not only degrades classification accuracy but also results in less reliable confidence estimates. Among the evaluated architectures, PTv3 consistently achieves the lowest ECE across all noise configurations, suggesting superior calibration performance compared to both PointNet and DGCNN.
Table 3: Expected calibration error for all models across noise levels. Lower is better.
Model None Light Moderate Heavy
PointNet 0.0561 0.0821 0.0848 0.0994
DGCNN 0.0493 0.0522 0.0553 0.0591
PTv3 0.0388 0.0223 0.0314 0.0419
[…]
To further investigate the relationship between input noise and calibration, we stratify the test samples into quartiles based on their mean measurement uncertainty $\sigma$ and report the Expected Calibration Error. For each model, we collect the values of $\sigma$ from all noise levels and divide them into four bins of equal size, ranging from the lowest to the highest uncertainty. Results are presented in Table 4.
For all models, ECE increases monotonically from Bin 0 (lowest uncertainty) to Bin 3 (highest uncertainty), confirming that measurement noise negatively impacts model calibration. While DGCNN maintains relatively stable ECE across bins, PointNet shows a more pronounced degradation in calibration with increasing noise. PTv3 again demonstrates the lowest ECE in all quartiles, indicating not only strong overall calibration but also robustness to varying levels of measurement uncertainty.
Table 4: Stratified expected calibration error computed across quartiles of measurement uncertainty $\sigma$. Bin 0 corresponds to the lowest $\sigma$ quartile and Bin 3 to the highest. Lower is better.
Model Bin 0 Bin 1 Bin 2 Bin 3
PointNet 0.0561 0.0799 0.0850 0.1018
DGCNN 0.0499 0.0512 0.0542 0.0616
PTv3 0.0395 0.0247 0.0308 0.0437
We plot calibration curves for each subset in Figure 3. Across all models, we observe a clear degradation in calibration as $\sigma$ increases. In the lowest quartile (Q1), models are relatively well-calibrated, with accuracy closely tracking confidence. However, from Q2 onward, calibration quality deteriorates, particularly in the low-confidence bins, where models tend to be significantly overconfident. This trend becomes more pronounced in Q3 and Q4, where curves increasingly deviate from the ideal diagonal. PTv3 maintains a modest advantage in calibration across most quartiles, while PointNet shows greater overconfidence in higher-noise regimes. These results align with the overall ECE trends and reinforce the sensitivity of model calibration to input-level measurement noise.
[…]
Model Light Moderate Heavy
PointNet 0.0726 0.0379 0.0437
DGCNN 0.1346 0.0739 0.0716
PTv3 0.1506 0.1697 0.1453
Correlation Between Measurement Uncertainty and Predicted Uncertainty
Table 5 reports the Pearson correlation between the ground-truth measurement uncertainty $\sigma$ and the predicted uncertainty (see Section 2).
Among the evaluated models, PTv3 consistently achieves the highest correlation across all noise levels, with values close to 0.19, suggesting that its confidence estimates are strongly aligned with the true measurement uncertainty. This indicates that PTv3 not only maintains a low Expected Calibration Error but also exhibits well-calibrated uncertainty estimates that appropriately respond to increases in noise level.
DGCNN also shows moderate positive correlation, achieving values between 0.07 and 0.13. While this is substantially higher than PointNet in all cases, the correlation decreases notably under moderate and heavy noise levels, indicating some limitations in DGCNN’s uncertainty awareness in more challenging conditions.
In contrast, PointNet demonstrates considerably lower correlation, with values near 0.04 under moderate and heavy noise, indicating that its predicted confidence is largely insensitive to the actual measurement uncertainty. These findings highlight that while all models can achieve high classification accuracy, their ability to recognize when predictions should be uncertain varies significantly. The strong performance of PTv3 suggests that modern transformer-based architectures may be better suited to capture and reflect input uncertainty, whereas simpler point-based networks are more prone to overconfident predictions regardless of noise level.
Training Dynamics for PTv3
We investigate the classification accuracy, calibration (ECE), and correlation between predicted uncertainty and measurement uncertainty for PTv3 across several training times and noise levels. For each noise level, we save the best model (in terms of classification accuracy) after 20, 40, 60, 80, and 100 training epochs.
Figure 4 shows our results. As expected, classification accuracy (left) improves steadily over time. However, ECE (center) tends to worsen as training progresses, especially under moderate and heavy noise. This suggests that while the model becomes more accurate, its confidence estimates become less aligned with actual correctness, indicating degraded calibration.
Most notably, the Pearson correlation between predicted and measurement uncertainty (right) also decreases with additional training epochs. This decline implies that the model becomes less sensitive to input degradation, even as it becomes more confident and accurate. Such a trend aligns with prior observations in the literature, where overfitting can lead to overconfident predictions and reduced uncertainty awareness (Guo et al. 2017).
These findings underscore the importance of evaluating not only accuracy but also calibration and uncertainty sensitivity when training deep learning models, especially for safety-critical applications. Rather than relying solely on classification accuracy as a stopping criterion, monitoring calibration and uncertainty awareness could provide more robust indicators for determining when to halt training.
Conclusion
We present ModelNet40-E, a novel uncertainty-aware benchmark designed to evaluate the robustness, calibration, and uncertainty sensitivity of point cloud classification models. We introduce LiDAR-like noise to the original ModelNet40 dataset at three distinct noise severity levels.
We evaluate three representative architectures—PointNet, DGCNN, and Point Transformer v3 (PTv3)—across classification accuracy, Expected Calibration Error (ECE), and uncertainty awareness (correlation between measurement and predicted uncertainty). Our results reveal that classification accuracy alone is insufficient to capture model reliability under corruption. For instance, while DGCNN achieves the highest accuracy across noise levels, PTv3 consistently demonstrates superior calibration and uncertainty sensitivity. Notably, PTv3 is the only model to maintain a relatively stable Pearson correlation between measurement uncertainty and predicted uncertainty as noise severity increases, reflecting a more robust and reliable uncertainty estimation. Moreover, our analysis of training dynamics shows that prolonged training can degrade calibration and uncertainty sensitivity even as accuracy continues to improve, suggesting the importance of monitoring these metrics as potential early stopping criteria.